# On the existence of weighted-cscK metrics

### v2

@@ 4.B. Closedness
    $\frac{\mathrm{tr}_{\mathrm{v},\omega}\theta}{\mathrm{w}(m_{\omega})}=\underline{\theta}.$
@@
$\sigma_{i}^{*}(\omega_{\varphi_{i}}^{n})=(\sigma_{i}^{*}\omega_{\varphi_{i}})^{n}=\omega_{\phi_{i}}^{n},\quad\sigma_{i}^{*}(\mathrm{v}(m_{\varphi_{i}})\omega_{\varphi_{i}}^{n})=\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}$
@@
[…]
@@
[…]
@@
Let $f_{i}=\frac{1-t_{i}}{t_{i}}u_{i}$. Let $\alpha=\alpha(X,\omega)$ be the alpha-invariant of the Kähler class$[\omega]$. Then there exists $C>0$ such that $\int_{X}e^{-p\cdot f_{i}}\omega^{n}<C$ for any $p<\frac{t_{i}\alpha}{1-t_{i}}$. Moreover, we have $e^{-f_{i}}\rightarrow 1$ in $L^{p}(\omega^{n})$ as $t_{i}\rightarrow 1$ for any $p<\infty$.
@@
$\phi_{i}\rightarrow\phi_{*}\text{ in }C^{1,\alpha}\text{ for any }0<\alpha<1\text{ and }\sqrt{-1}\partial\overline{\partial}\phi_{i}\rightarrow\sqrt{-1}\partial\overline{\partial}\phi_{*}\text{ weakly in }L^{p},$ (4.12)
@@
$\qquad\qquad\qquad\qquad\nabla(\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i})\rightarrow\nabla F_{*}\text{ weakly in }L^{p}.$
@@
$\omega_{\phi_{i}}^{k}\rightarrow\omega_{\phi_{i}}^{k}\text{ weakly in }L^{p}\text{ for any }1\leq k\leq n\text{ and }p<\infty.$
@@
To obtain the second equation, by (2.10), we first re-write the equation (4.8) to
@@
For the right hand side, by (4.13), (4.12), $\phi_{i}\rightarrow\phi_{*}$ in $C^{1,\alpha}$ for any $10<\alpha<$ 1. Then following the construction in the proof of [HL23, Proposition 10], we have
@@
In the proof of Lemma 4.11, we have obtain that $\phi_{*}\in W^{3,p}$ for any $p<\infty$. Hence $\omega_{\phi_{*}}\in C^{\alpha}$ for any $10<\alpha<$ 1. By (4.14) and Schauder estimate, we derive $F_{*}\in C^{2,\alpha}$ for any $10<\alpha<$ 1. Then the higher regularity follows from bootstrap (see Corollary 5.3). ∎
@@ 5.A. Apriori estimates for weighted cscK equation
$\mathrm{Bisec}(\omega)\geq-B,\quad\mathrm{Ric}(\omega)\leq A_{1}\omega,\quad\|F\|_{L^{\infty}}+\|\phi\|_{L^{\infty}}\leq C_{0}.$
@@
By Schauder estimate, we deduce $\|\phi\|_{C^{3,\alpha^{\prime}}}\leq C_{4}$. Going back to (2.13b) again, the coefficient are in $C^{\gamma}$ for any $10<\gamma<$ 1, hence we have $\|F\|_{C^{1,\gamma}}\leq C_{5}$. Now, the coefficients of (5.8) are in $C^{\gamma}$ for any $10<\gamma<$ 1. Thus we derive $\|\phi\|_{C^{3,\gamma}}\leq C_{6}$ for any $10<\gamma<$ 1.
@@ 5.B. Linearization of opeartors
where $-\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big{(}J\xi^{\alpha}(\psi)\big{)},\overline{\partial}\dot{\phi}\rangle_{\phi}=\langle\theta,\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi}\rangle$.
@@
The linearized operator $\frac{d}{dt}{\mathcal{T}}-\langle\partial{\mathcal{T}},\overline{\partial}\cdot\rangle_{\phi}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$- self-adjoint second order elliptic linear operator.

### v3

@@ 2. $(2)$
    $\frac{\mathrm{tr}_{\mathrm{v},\omega}\theta}{\mathrm{w}(m_{\omega})}=\underline{\theta}.$ (4.8)
    Indeed, the equation (4.8) can be rewritten as
    $\frac{\mathrm{d}^{*}(\mathrm{v}(m_{\omega})\mathrm{d}\psi_{0})}{\mathrm{w}(m_{\omega})}=\frac{\mathrm{tr}_{\mathrm{v},\omega}(\omega)}{\mathrm{w}(m_{\omega})}-\underline{\theta}.$
    Note that $\frac{\mathrm{d}^{*}(\mathrm{v}(m_{\omega})\mathrm{d}\psi_{0})}{\mathrm{w}(m_{\omega})}$ a second order $(\cdot,\cdot)_{\mathrm{w},\omega}$-self-adjoint elliptic operator. By standard elliptic theory it follows that there exist a solution $\psi_{0}\in C^{\infty}(X)_{K}$ of (4.8).
@@
$\sigma_{i}^{*}(\omega_{\varphi_{i}}^{n})=(\sigma_{i}^{*}\omega_{\varphi_{i}})^{n}=\omega_{\phi_{i}}^{n},\qquad\sigma_{i}^{*}(\mathrm{v}(m_{\varphi_{i}})\omega_{\varphi_{i}}^{n})=\mathrm{v}(m_{\phi_{i}})\omega_{\phi_{i}}^{n}$
@@
$\text{L.H.S.}=\sigma_{i}^{*}\left(\left(\Delta_{\varphi_{i}}-\frac{\mathrm{v}_{,\alpha}(m_{\varphi_{i}})}{\mathrm{v}(m_{\varphi_{i}})}J\xi^{\alpha}\right)F_{i}\right)=\left(\Delta_{\phi_{i}}-\frac{\mathrm{v}_{,\alpha}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}J\xi^{\alpha}\right)F_{i}\circ\sigma_{i},$ (4.14)
@@
$\text{R.H.S.}=-\ell_{\mathrm{ext}}(m_{\phi_{i}})\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+(1-\frac{1}{t})\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}(\sigma_{i}^{*}\theta)}{\mathrm{v}(m_{\phi_{i}})}-(1-\frac{1}{t})\underline{\theta}\frac{\mathrm{w}(m_{\phi_{i}})}{\mathrm{v}(m_{\phi_{i}})}+\frac{\mathrm{tr}_{\mathrm{v},\phi_{i}}\mathrm{Ric}(\sigma_{i}^{*}\omega)}{\mathrm{v}(m_{\phi_{i}})}.$ (4.15)
@@
[…]
@@
[…]
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt\nabla(\widetilde{F}_{i}-\log\mathrm{v}(m_{\phi_{i}})+f_{i})\rightarrow\nabla F_{*}\text{ weakly in }L^{p}.$
@@
[…]
@@
To obtain the second equation, by (2.B), we first re-write the equation (4.13) to
@@
For the right hand side, by (4.18), (4.17), $\phi_{i}\rightarrow\phi_{*}$ in $C^{1,\alpha}$ for any $0<\alpha<1$. Then following the construction in the proof of [HL20, Proposition 10], we have
@@
In the proof of Lemma 4.11, we have obtain that $\phi_{*}\in W^{3,p}$ for any $p<\infty$. Hence $\omega_{\phi_{*}}\in C^{\alpha}$ for any $0<\alpha<1$. By (4.19) and Schauder estimate, we derive $F_{*}\in C^{2,\alpha}$ for any $0<\alpha<1$. Then the higher regularity follows from bootstrap (see Corollary 5.3). ∎
@@ 5.A. Apriori estimates for weighted cscK equation
$\mathrm{Bisec}(\omega)\geq-B,\qquad\mathrm{Ric}(\omega)\leq A_{1}\omega,\qquad\|F\|_{L^{\infty}}+\|\phi\|_{L^{\infty}}\leq C_{0}.$
@@
By Schauder estimate, we deduce $\|\phi\|_{C^{3,\alpha^{\prime}}}\leq C_{4}$. Going back to (2.14b) again, the coefficient are in $C^{\gamma}$ for any $0<\gamma<1$, hence we have $\|F\|_{C^{1,\gamma}}\leq C_{5}$. Now, the coefficients of (5.8) are in $C^{\gamma}$ for any $0<\gamma<1$. Thus we derive $\|\phi\|_{C^{3,\gamma}}\leq C_{6}$ for any $0<\gamma<1$.
@@ 5.B. Linearization of operators
where $\sum_{\alpha}\mathrm{v}_{,\alpha}\langle\partial\big(J\xi^{\alpha}(\psi)\big),\overline{\partial}\dot{\phi}\rangle_{\phi}=\langle\theta,\partial\mathrm{v}\wedge\overline{\partial}\dot{\phi}\rangle$.
@@
The linearized operator $\frac{d}{dt}{\mathcal{T}}-\langle\partial{\mathcal{T}},\overline{\partial}\cdot\rangle_{\phi}$ is a $(\cdot,\cdot)_{\mathrm{w},\phi}$- self-adjoint second order elliptic linear operator. In particular, if $\theta$ is positive, then the operator has negative first eigenvalue for $R>0$.